# MCiteBench: A Benchmark for Multimodal Citation Text Generation in MLLMs

### v2

@@ 5.1 Evaluation Settings
In this work, citations for textual modalities are indicated using box brackets (e.g., [1][2]), while for figures and tables, we adopt the index provided in their captions (e.g., Figure 3, Table 2). For both single-source and multi-source evidence questions, the multimodal corpus $M$ comprises 5 items, including the ground truth evidence and distractors. Distractors are randomly selected from other content within the same paper.
@@
In this study, we use GPT-4o to assess the entailment relationship between model responses and their cited evidence  [^4] We validate GPT-4o’s reliability in Appendix B.3..
@@
Among the open-source models, we choose InternVL-2.5 (8B and 78B) Chen et al. (2024), Qwen2-VL (7B and 78B) Wang et al. (2024b), Llama 3.2-Vision (11B and 90B) Meta (2024), as well as Llava-OneVision (and its chat version) and MiniCPM-V-2.6 Yao et al. (2024). For proprietary models, we test GPT-4o (GPT-4o-2024-11-20) and GPT-4o-mini (GPT-4o-mini-2024-07-18) Hurst et al. (2024).
@@ 5.2 Main Results
Model performance reflects the difficulty of the questions, with higher accuracy scores observed on Locating questions compared to Explanation questions, indicating that Explanation questions are more challenging. As shown in Table 4, as question difficulty increases, model citation performance tends to decrease. For instance, GPT-4o achieves 85.74% in Source F1 for single source Locating questions but drops to 56.82% for single source Explanation questions. Explanation questions place higher demands on citation generation, as they require an in-depth analysis of the inputs.
Can MLLMs Cite Accurately in Multi-Source Scenarios?
In multi-source scenarios, models achieve higher Citation F1 and Source F1 scores, as the ground truth citations contain more items. However, the stricter Source Exact Match metric is lower than in single-source scenarios. This highlights the challenge of citing in multi-source scenarios, where models must correctly include relevant sources while avoiding irrelevant ones.
Figure 5: Results of MCiteBench across different modalities. Source Exact Match(S-EM) score indicates whether the model’s response-level citation is the same as the ground truth.
@@ 5.3 Analysis
In this section, we discuss several research questions, revealing the inherent biases and bottlenecks in the task.
RQ1: Do Models Prefer Specific Modalities for Citation?
@@
Figure 6: Attention distribution across multimodal sources. Visual sources include figures and tables.
RQ2: How Do Models Allocate Attention When Generating Source Citations?
Figure 7: Attention heatmap for ground truth and distractors. The heatmap shows the model’s attention distribution when predicting the next token of “Logic-LM has higher accuracy. According to Figure $\Diamond$” in its response. Although the model answers correctly, its attention in the distractors remains focused on index positions (e.g., [1], Figure 2).
@@
We analyze the attention distribution of Qwen2-VL-7B when generating such tokens. Specifically, we focus on its behavior when predicting the next token after “Logic-LM has higher accuracy. According to Figure $\Diamond$ in its response. Notably, the distractors are sampled from unrelated papers, meaning they provide no useful information for answering the question.
@@ Results
As shown in Figure 7, the model’s attention heatmap reveals an intriguing pattern: even when the response is based entirely on a specific piece of evidence, the model’s attention is not solely focused on it. When generating the token after “According to Figure”, the model’s attention remains high on textual index positions (e.g., “[1]”, “[2]”), even though the context suggests the model should focus on figure evidence. This suggests that while the model correctly cites the source, it maintains a broader contextual awareness by attending to multiple potential evidence.
Model Understanding Attribution
Open-Source(7-14B)
Qwen2-VL-7B 0.94 0.59
InternVL2_5-8B 0.91 0.56
Open-Source(>70B)
Qwen2-VL-72B 0.95 0.62
InternVL2_5-78B 0.93 0.62
Proprietary
GPT-4o-mini 0.93 0.61
GPT-4o 0.94 0.63
Table 3: Comparison of model performance on Understanding and Attribution tasks. Both tasks share the same input sources, but differ in the nature of the questions: one requires answering a question, while the other asks which source can help answer the question.
RQ3: Understanding or Attribution: What is the Bottleneck?
Insights from RQ2 show that even when models generate responses based on the correct evidence, they still focus heavily on distractors. This suggests a possible gap between understanding the content and correctly attributing it. This raises a critical question: Does the bottleneck in multimodal citation text generation stem from limitations in understanding or attribution?
On one hand, multimodal understanding requires the model to grasp the meaning of the inputs and use them to generate a coherent response. On the other hand, multimodal attribution demands that the model’s generated response can be traced back to specific, verifiable sources in the input.
Settings
We design multiple-choice questions for understanding and attribution. We constructed single source Locating QA pairs and sampled distractors from unrelated papers, ensuring that the questions remain straightforward to answer. Understanding questions involve a 4-option multiple-choice format, while attribution questions use a 5-option multiple-choice format, with each option corresponding to one of the five information sources in the input.
In the attribution task, models select a source identifier (e.g., “[1]”, “Figure 2”) from a fixed set of options, directly reflecting models internal computation during citation generation, where they select the source identifier tokens based on the probability of the next token.
Results
Results in Table 3 show that while models achieve over 90% accuracy on understanding questions, they perform worse on attribution questions. This suggests that the bottleneck lies not in multimodal understanding but in multimodal attribution. This finding underscores a fundamental limitation in current multimodal models: they can process and understand multimodal inputs well but struggle to attribute outputs to the correct evidence accurately.
@@ 6 Conclusion
In this paper, we introduce MCiteBench for evaluating multimodal citation text generation in MLLMs. We build MCiteBench using academic papers and their review-rebuttal interactions, ensuring a high-quality benchmark for evaluating multimodal citation text generation. Leveraging this benchmark, we conduct a detailed evaluation of model performance across multiple dimensions, including source reliability, answer accuracy, and citation quality. Through extensive experiments, we find that existing models struggle to accurately attribute their outputs to the correct multimodal sources. Furthermore, we dive deep into the analysis of attention distribution during citation generation and uncover biases and bottlenecks that hinder accurate attribution. We hope that MCiteBench provides valuable insights for citation text generation tasks and contributes to the development of models capable of generating faithful and verifiable responses.

### v3

@@ 5.1 Evaluation Settings
In this work, we indicate citations from textual content using box brackets (e.g., “[1]”), and refer to figures and tables by the indices in their captions (e.g., “Figure 3”, “Table 2”). We conduct an ablation study to assess the impact of including figure captions in the input. [^4] See Table 13 for details in Appendix C.2. For both single-source and multi-source evidence questions, the multimodal corpus $M$ comprises 5 items, including the ground truth evidence and distractors. Distractors are randomly selected from other content within the same paper.
@@
In this study, we use GPT-4o to assess the entailment relationship between model responses and their cited evidence. [^5] We validate GPT-4o’s reliability in Appendix B.3, and further verify in Appendix C.1 that it does not exhibit strong self-preference when evaluating responses in our task.
@@
For open-source models, we test InternVL-2.5 (8B/78B) Chen et al. (2024), Qwen2-VL (7B/78B) Wang et al. (2024b), Llama 3.2-Vision (11B/90B) Meta (2024), Llava-OneVision (and its chat version) and MiniCPM-V-2.6 Yao et al. (2024). For proprietary models, we test GPT-4o (GPT-4o-2024-11-20) and GPT-4o-mini (GPT-4o-mini-2024-07-18) Hurst et al. (2024).
@@ 5.2 Main Results
Model performance reflects the difficulty of the questions, with higher accuracy scores observed on locating questions compared to explanation questions, indicating that explanation tasks are more challenging. As shown in Table 4, as question difficulty increases, model citation performance tends to decrease. For instance, GPT-4o achieves 85.74% in Source F1 for single source locating questions but drops to 56.82% for single source explanation questions. Explanation questions place higher demands on citation generation, as they require an in-depth analysis of the inputs.
How Do Multi-Source Scenarios Affect Generating Text with Citations in MLLMs?
In multi-source settings, models tend to achieve higher Citation F1 and Source F1 scores, as multiple valid references allow for partial credit. Unlike single-source questions with only one correct citation, multi-source questions permit credit for correctly identifying any subset of the ground truth, naturally resulting in higher metric values. However, the stricter Source Exact Match metric is lower than in single-source scenarios. This highlights the challenge of citing in multi-source scenarios, where models must correctly include relevant sources while avoiding irrelevant ones.
@@ 5.3 Analysis
In this section, we discuss several research questions, revealing the inherent biases in the task.
RQ1: Can MLLMs Accurately Identify the Source Needed to Answer a Question?
Model Overall By Modality
Figure Table Text
Open-Source(7-14B)
Qwen2-VL-7B-Instruct 0.45 0.40 0.38 0.55
InternVL2_5-8B 0.48 0.37 0.42 0.65
Open-Source(>70B)
Qwen2-VL-72B-Instruct 0.59 0.50 0.57 0.71
InternVL2_5-78B 0.58 0.51 0.50 0.72
Proprietary
gpt-4o-mini 0.52 0.47 0.48 0.61
gpt-4o-2024-11-20 0.60 0.52 0.55 0.73
Table 3: Model accuracy on identifying the most relevant source for answering a question under the multi-choice setting.
Generating text with citation can be abstracted into a two-stage process: (1) generating a response, and (2) mapping that response to the appropriate supporting input sources by producing attribution tokens such as “[1]” or “Figure 3”.
Instead of requiring the model to generate an answer and then attribute it, we directly evaluate its ability to identify which source would be most helpful in answering a given question. Specifically, we ask: Can a model identify the correct source needed to answer a given question?
Settings
We construct a probing task based on Single-Source Explanation QA. For each example, we provide the model with a question and 5 candidate sources (1 correct + 4 distractors). The model is tasked with selecting which source would be most helpful in answering the question. [^6] See Table 10 for details in Appendix A.3.
Results
Results are presented in Table 3. Importantly, this task directly evaluates the model’s ability to identify relevant sources based solely on the question, rather than relying on model-generated answers or intermediate claims. Despite this seemingly simplified setting, no model achieves more than 60% accuracy, highlighting the persistent difficulty in accurately grounding questions in the correct source.
In addition, we observe a consistent performance gap across modalities: models perform better when reasoning over textual sources compared to visual inputs such as figures and tables, which leads to our next research question.
RQ2: Does Modality Influence Citation Performance?
Figure 5: Source Exact Match score of models on the MCiteBench benchmark across different modalities, under the multi-source explanation setting with two gold evidence items per question.
Figure 6: Attention distribution across multimodal sources.
@@ Results
Figure 7: Attention heatmap during source reference generation. The heatmap shows how the model distributes attention when generating the next token in its response, continuing the sentence “Logic-LM has higher accuracy. According to Figure $\Diamond$. Although the model answers correctly, its attention in the distractors remains focused on index positions (e.g., “[1]”, “Figure 2”).
@@
RQ3: What Do Models Look At When Generating Citations?
@@
Specifically, we examine the attention patterns of Qwen2-VL-7B when continuing a partially generated sentence ending in “According to Figure”, and tasked with predicting the next token (e.g., “4”). This allows us to assess which input regions the model attends to when making source attribution decisions.
Specifically, we focus on its behavior when predicting the next token after “According to Figure $\Diamond$ in its response. Notably, the distractors are sampled from unrelated papers, meaning they provide no useful information for answering the question.
@@
As shown in Figure 7, the model’s attention heatmap reveals an intriguing pattern: even when the response is based entirely on a specific piece of evidence, the model’s attention does not solely focus on it. When generating the token after “According to Figure”, the model’s attention remains high on textual index positions (e.g., “[1]”, “[2]”), even though the context suggests the model should focus on figure evidence. This suggests that while the model correctly cites the source, it maintains a broader contextual awareness by attending to multiple potential evidence.
@@ 6 Conclusion
In this paper, we introduce MCiteBench, a high-quality benchmark built from academic papers and their review–rebuttal interactions, to evaluate the ability of MLLMs to generate text with citations from multimodal input. Leveraging this benchmark, we conduct a detailed evaluation of model performance across multiple dimensions. Through extensive experiments, we find that existing models struggle to accurately attribute their outputs to the correct multimodal sources. Furthermore, we dive deep into the analysis of attention distribution during citation generation and uncover modality bias exhibited by current models. We hope that MCiteBench offers valuable insights into generating text with citations and contributes to the development of models capable of producing faithful and verifiable responses.
@@ 2. Your scoring should be strict, and follow these guidelines:
A.3 Source Identification Prompt
We list the prompt used to evaluate whether a model can identify the most relevant source for answering a given question in Table 10.
Please identify the most relevant source of evidence to locate information that could address the following query. Provide your answer by selecting one of the options: A, B, C, D, or E. Begin your response with the selected letter and, if necessary, briefly explain why it is the most relevant source.
Identify where information about
{question}
can be found.
Which source is most relevant?
{options}
Table 10: Prompt used to evaluate source identification ability in the multi-choice setting.
@@ B.2 Quality Control
Our annotation process involves three students from the artificial intelligence field, with one serving as the annotation lead. The process takes approximately one month to complete, and annotators are compensated at the local minimum hourly wage rate. Regarding inter-annotator agreement, in cases of disagreement about whether to retain specific data points, the annotation lead makes the final decision.
@@ B.3 Agreement Between Human Annotations and GPT-4o
Appendix C Ablation Study
C.1 Effect of LLM Judge Choice
To further investigate the robustness of our automatic evaluation setup, we conduct an ablation study using alternative judge models. Specifically, we evaluate model performance on 90 randomly sampled examples (30 Locating and 60 Explanation questions), comparing scores assigned by GPT-4o and DeepSeek V3-0324.
As shown in Table 12, GPT-4o consistently achieves the highest scores under both judge models. While DeepSeek V3-0324 tends to yield slightly higher absolute scores across all models, the relative ranking remains consistent. This suggests that self-preference bias from GPT-4o does not significantly affect evaluation outcomes, confirming the robustness of our LLM-based evaluation setup.
Model GPT-4o Judge DeepSeek V3 Judge
Proprietary
GPT-4o-2024-11-20 66.72 76.11
GPT-4o-mini 64.55 72.22
Open-Source(7-14B)
Qwen2-VL-7B-Instruct 63.77 65.00
InternVL2_5-8B 62.53 67.78
Table 12: Accuracy of citation evaluation across models under different LLM judges on a 90-sample subset.
C.2 Effect of Captions on Citation Performance
To assess the role of visual-textual information in multimodal citation understanding, we conduct a comprehensive ablation study across the full benchmark dataset (3,000 examples), comparing model performance with and without image captions.
As shown in Table 13, we observe that including captions leads to minor changes in performance across most evaluation metrics. Notably, the accuracy improvements are modest for both GPT-4o-mini and GPT-4o. Interestingly, in some cases (e.g., GPT-4o-mini), the inclusion of captions slightly degrades performance in label prediction and citation generation (as measured by F1 and exact match), while GPT-4o exhibits a substantial gain in citation F1.
These results demonstrate that our benchmark does not rely solely on OCR-extracted text, and that the image-caption setting we adopt provides a reasonable and realistic testbed for evaluating MLLMs’ citation capabilities. At the same time, the relatively limited gains from caption inclusion highlight that current models still face challenges in grounding their responses effectively, even when textual cues are explicitly embedded in the image.
Model GPT-4o-mini GPT-4o
No Cap. Acc 62.25 64.92
With Cap. Acc 64.55 66.72
Acc Impr. +2.30 +1.80
No Cap. S-F1 48.63 67.10
With Cap. S-F1 45.34 68.19
S-F1 Impr. -3.29 +1.09
No Cap. S-EM 28.98 38.47
With Cap. S-EM 25.34 39.11
S-EM Impr. -3.64 +0.64
No Cap. C-F1 55.51 76.54
With Cap. C-F1 49.53 87.43
C-F1 Impr. -5.98 +10.89
Table 13: Effect of captions on citation evaluation performance across multiple metrics. The performance with captions is compared to that without captions.